# Combinatorial Auctions without a Numeraire: The Case of Blockchain Trade-Intent Auctions

### v1

@@ 2.1 Timing and mechanism
    Then, each solver $i\in\{1,2\}$ sends three bids. Two bids are bids on individual orders, denoted by $q_{i}(B)$ and $q_{i}(D)$, in units of asset $B$ and $D$, respectively. The third bid is a tuple $Q_{i}=\{Q_{i}(B),Q_{i}(D)\}$, interpreted as a bid for the entire batch, that is, a bid for both orders, where $Q_{i}(B)$ is in units of asset $B$ and $Q_{i}(D)$ is in units of asset $D$.
@@
An important observation is that the feasibility constraint also has implications for the set of feasible mechanisms. Intuitively, a mechanism is not feasible if, for all non-trivial bids by a solver $i\in\{1,2\}$, there are inventories for solver $i$ and bids by the other solvers such that solver $i$ violates the feasibility constraint. Conversely, a mechanism is feasible if, for any possible inventory, solvers can make non-trivial bids and be guaranteed to respect the feasibility constraint.
@@
A mechanism $\Gamma$ is feasible if $\forall i\{1,2\}$,
$\exists\{q_{i}(B),q_{i}(D),Q_{i}(B),Q_{i}(D)\}\geq\{\underline{\beta},\underline{\delta},\underline{\beta},\underline{\delta}\}$
with at least one strict inequality, such that
@@ 3 Batch auction and simultaneous standard auctions
We can use the above framework to study the two dominant mechanisms for trade intent auctions: simultaneous standard auctions (as implemented by 1inch Fusion and UniswapX in the form of Dutch auctions) and batch auctions (as implemented by CoW Swap). These mechanisms are quantitatively important: in May 2024, 1-inch fusion and UniswapX intermediated USD 650 million and USD 1.6 billion, respectively, while CoW Swap intermediated USD 2.4 billion. These mechanisms are also relevant theoretical benchmarks because they are sub-cases of the above model: simultaneous standard auctions are equivalent to the above mechanism in which the batch bids are ignored, while batched auctions are equivalent to the above mechanism in which individual bids are ignored.
@@ 3.1 Simultaneous standard auctions
If there are two simultaneous standard auctions, then in equilibrium, solver 1 wins the first order and returns $\underline{\beta}$ units of asset $B$, while solver 2 wins the second order and returns $\underline{\delta}$ units of asset $D$ [^6] Note that, technically, there could be other equilibria because the least productive solver on each trade never wins and hence could bid any amount that satisfies its feasibility constraint. If the auction is, for example, in second price, then some equilibria are even worse for the traders than that described in the Lemma. These equilibria are, however, not robust to perturbations of the type space. For example, if solver 1 can produce $\underline{\delta}+\epsilon$ for $\epsilon\rightarrow 0$ when it wins only trade 2, and solver 2 can produce $\underline{\beta}+\epsilon$ when it wins only trade 1, then there an infinitesimally small probability that solver 1 wins trade 2 and solver 2 wins trade 1. Then, the equilibrium described in the Lemma becomes unique.
@@ 3.2 Batch auction
To start, remember that feasibility prevents “second price” mechanisms that use the losing bid to determine the assets returned by the winning solver: even if the winning bid has a higher value, it may nonetheless deliver less of one asset than the second-highest bid. We, therefore, restrict our attention to a “first price” (or “pay your bid”) mechanism in which the winning solver returns the assets specified in its batched bid, which is the approach followed by CoW Swap. [^8] In practice, the batch auction model implemented by CoW Swap also has a subsidy to solvers, which we ignore for simplicity.
@@
The important observation is that, in the batch auction, the total value of the assets returned to the traders always exceeds that of the simultaneous individual-trade auctions. The reason is that batching forces the two solvers to compete: even if there is a single “strong” solver per trade (and little competition if trades are auctioned off independently), there are two “strong” solvers in the batch auctions, and hence higher competition between solvers. At the same time, there is an indeterminacy relative to what each trader receives. There are equilibria in which a trader receives less in a batch auction than what this trader would have received in the simultaneous individual-trade auctions. We summarize these observations in the following remark.
@@
There are equilibria of the batch auction in which traders receive more than $\underline{\beta}$ and $\underline{\delta}$, and both traders are better off relative to the simultaneous standard auctions (c.f, Lemma 1). There are also equilibria in which either trader 1 receives less than $\underline{\beta}$ or trader 2 receives less than $\underline{\delta}$, that is, there are equilibria in which one of the two traders receives less in a batch auction than in the simultaneous standard auctions.
@@
The goal of this section is to formalize a notion of fairness applicable to a generic combinatorial auction $\Gamma$. Doing so requires establishing a counterfactual mechanism against which to compare the outcome of $\Gamma$. The key observation is that such counterfactual is part of the mechanism $\Gamma$, because it depends on the bids and determines the equilibrium allocation and payments. To say it differently, the reference for fairness must be generated endogenously as part of the mechanism.
@@
The first part of the above assumption imposes a mild form of monotonicity in the auction’s allocation. The second part is a form of independence of the auction’s payments whenever the individual bids determine the outcome. It is slightly more restrictive than the first part but still quite mild, especially in conjunction with the requirement that the mechanism is feasible.
Remember that feasibility requires that a solver who does not win trade 1 does not need to return any asset B and, similarly, a solver who does not win trade 2 does not need to return any asset D. It also implies that if a solver wins a trade, there is an upper bound to the assets that such a solver needs to return. This upper bound is solely a function of the winner’s bids so that, when bidding, a solver can be sure not to violate the feasibility constraint for any bids by the opponent. If, for simplicity, we interpret this upper bound as the bid on the individual trade itself, then feasibility implies that the winner of trade 1 (or 2) returns an amount of assets B (or D) that is between zero and its bid on trade 1 (or 2). The above assumption then further implies that, in case the outcome is determined by the individual bids, where exactly the winner’s payment lies in the feasible interval is independent of the batched bids.
@@ 5 Equilibrium of the fair combinatorial auction
The remainder of the section, therefore, derives the equilibrium of different types of fair combinatorial auctions. For ease of comparison with the above benchmarks, we restrict our attention to mechanisms such that:
@@
    If multiple batched bids are fair, then the winner is the one with higher market value. [^9] This assumption is both for simplicity and for consistency with the standard batch auction. Also, the intuition behind our results extends to other monotonically increasing functions of the assets returned to the users.
@@
Suppose that the individual, item-by-item auctions are second-price auctions. Formally, define the solver with the best batched-bid as $W\equiv\mbox{argmax}_{i\in\{1,2\}}\left\{Q_{i}(B)+p_{DB}Q_{i}(D)\right\}$. Then, the mechanism is:
  • if $Q_{W}(B)>\min\{q_{1}(B),q_{2}(B)\}$ and $Q_{W}(D)>\min\{q_{1}(D),q_{2}(D)\}$, then $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$, and all other transfers are zero. [^10] Also here, there are some technicalities related to how ties are broken in case a solver matches the opponent’s batched bid with its individual-trade bid. The formulation in the text assumes that, in case of ties, the batched bid is disqualified as unfair. This assumption guarantees the existence of the equilibrium; otherwise, a solver may want to outbid the opponent’s batched bid by the smallest possible amount, but because the action space is continuous, this “smallest possible amount” is not defined. We also note that an alternative to imposing such a tie-breaking rule is to introduce an endogenous tie-breaking rule as in Jackson et al. (2002), which also guarantees that the equilibrium when the action space is discrete converges to the equilibrium with continuous action space as the action space becomes finer (see also the discussion in footnote7).
  • Otherwise, the best bid on a given order wins it and pays the second-highest bid. Formally, denote the winner of order 1 by $w_{1}\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(B)\}$, and the winner of order 2 by $w_{2}\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(D)\}$. The mechanism in this case is: $f(m)=\{w_{1},w_{2}\}$, $b_{w_{1}}(m)=\min\{q_{1}(B),q_{2}(B)\}$, $d_{w_{2}}(m)=\min\{q_{1}(D),q_{2}(D)\}$ and all other transfers are zero.
Consider any strictly positive $Q_{1}(B),Q_{2}(B),Q_{1}(D),Q_{2}(D)$ that are an equilibrium of the batch auction discussed earlier. The key observation is that these bidding strategies are also an equilibrium of the fair combinatorial auction with individual bids in second price whenever $q_{1}(B)=q_{2}(B)<\min\{Q_{1}(B),Q_{2}(B)\}$ and $q_{1}(D)=q_{2}(D)<\min\{Q_{1}(D),Q_{2}(D)\}$. The reason is that no individual solver can manipulate the reference for fairness. That is, it is possible that, in equilibrium, $q_{1}(B)=q_{2}(B)<Q_{i}(B)<\underline{\beta}$ $\forall i\{1,2\}$; nonetheless, no solver can change its individual-trade bid and disqualify the opponent’s bid as unfair. Finally, the above argument holds for any $Q_{1}(B),Q_{2}(B),Q_{1}(D),Q_{2}(D)$ strictly positive, while it fails (due to our tie-breaking rule) whenever the winning bid of the batch auction delivers zero assets to a trader, leading to the following lemma.
@@
The set of equilibria of the fair combinatorial auction with second-price individual trade auctions (as in 2) is the set of equilibria of the batch auction in which each trader receives strictly positive amounts.
Intuitively, when the individual-trade auctions are in second price, each individual solver has limited ability to manipulate the reference for fairness. Depending on the players’ bids, it is even possible that the outcome of the auction gives almost no asset to one player, and nonetheless, no solver can disqualify such an outcome as unfair. As a consequence, this auction reduces to, again, the batch auction (c.f., Lemma 2), with the only difference that delivering exactly zero assets to a trader is not an equilibrium here (but it is an equilibrium of the batch auction).
@@
The mechanism, in this case, is:
  • if $Q_{W}(B)>\max\{q_{1}(B),q_{2}(B)\}$ and $Q_{W}(D)>\max\{q_{1}(D),q_{2}(D)\}$, then $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$, and all other transfers are zero.
  • Otherwise, $f(m)=\{w_{1},w_{2}\}$, $b_{w_{1}}(m)=q_{w_{1}}(B)$, $d_{w_{2}}(m)=q_{w_{2}}(D)$ and all other transfers are zero.
We start with a useful preliminary result.
@@
Suppose that, in equilibrium:
$\begin{split}Q^{*}_{1}(B)+p_{DB}Q^{*}_{2}(D)\geq g\underline{\beta}+p_{DB}g\underline{\delta}Q^{*}_{2}(B)+p_{DB}Q^{*}_{2}(D)\geq g\underline{\beta}+p_{DB}g\underline{\delta},\end{split}$ (4)
that is, each solver’s batched bid returns at least the minimum value produced. Then, it must be that $Q^{*}_{1}(D)=g\cdot\underline{\delta},\leavevmode\nobreak\ Q^{*}_{1}(B)\in[g\underline{\beta},g\beta)$, and similarly $Q^{*}_{2}(B)=g\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)\in[g\underline{\delta},g\cdot\delta)$.
The lemma shows that, in its batched bid, each solver bids the maximum on the trade in which it has a disadvantage. The reason is that a solver’s batched bid is more likely to be considered unfair due to the amount of assets returned on the “weak” order (which is the opponent’s “strong” order). On the other component of the batched bid, each solver delivers more than the minimum amount but less than the maximum amount.
@@
$\begin{cases}(\beta-q_{1}(B))\mbox{pr}\left\{q_{2}(D)\geq g\underline{\delta}\right\}+\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)\geq g\underline{\beta}\\ \left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}&\mbox{ otherwise }\end{cases}$
where, again, $V\equiv g\underline{\beta}-p_{DB}\cdot g\underline{\delta}$. Note that, as long as $q_{1}(B)\geq g\underline{\beta}$, whether solver 1 wins one or both trades is independent of its bids, and solver 1 should set $q_{1}(B)=Q_{1}(B)=g\underline{\beta}$. Using this, we can rewrite solver 1’s payoff as
$\begin{cases}(\beta-g\underline{\beta})\mbox{pr}\left\{q_{2}(D)\geq g\underline{\delta}\right\}+\left(\beta\cdot g-g\underline{\beta}\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)=g\underline{\beta}\\ \left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}&\mbox{ otherwise }\end{cases}$
A second observation is that, as long as (4) holds, then $Q_{1}(B)\geq g\underline{\beta}$ and
$\left(\beta\cdot g-g\underline{\beta}\right)$ $\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}>$
$\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}$
and solver 1 always prefers $q_{1}(B)=g\underline{\beta}$ (i.e., kicking out the opponent’s batched bid) to $q_{1}(B)<g\underline{\beta}$ (i.e., not kicking out the opponent’s batched bid). However, $q_{1}(B)=g\underline{\beta}$ may not be feasible. Finally, note that when $q_{1}(B)<g\underline{\beta}$, then solver 1 payoff does not depend on $q_{1}(B)$ and, without loss of generality, we can assume $q_{1}(B)=\underline{\beta}$.
It follows that solver 1 best response is:
$\begin{cases}q_{1}(B)=Q_{1}(B)=g\underline{\beta}&\mbox{ if }\beta\geq g\underline{\beta}\\ q_{1}(B)=\underline{\beta},\leavevmode\nobreak\ Q_{1}(B)=\tilde{Q}_{1}(B)&\mbox{ otherwise }\end{cases}$
@@
$\tilde{Q}_{1}(B)\equiv\mbox{argmax}_{Q_{1}(B)}\left\{\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}\right\}$
Repeating the same reasoning for solver 2 yields the best response:
$\begin{cases}q_{2}(D)=Q_{2}(D)=g\underline{\delta}&\mbox{ if }\delta\geq g\underline{\delta}\\ q_{2}(D)=\underline{\delta},\leavevmode\nobreak\ Q_{2}(D)=\tilde{Q}_{2}(D)&\mbox{ otherwise }\end{cases}$
@@
$\tilde{Q}_{2}(D)\equiv\mbox{argmax}_{Q_{2}(D)}\left\{p_{DB}\left(\delta\cdot g-Q_{2}(D)\right)\mbox{pr}\left\{p_{DB}Q_{2}(D)+V>Q_{1}(B)|q_{1}(B)<g\underline{\beta}\right\}\right\}$
@@
Figure 1: The two best responses
@@
Suppose the individual trade auctions are in first price as in (1). Then, in equilibrium, (4) holds and:

### v2

@@ 2.1 Timing and mechanism
    Then, each solver $i\in\{1,2\}$ sends three bids. The first two are on individual orders, and we denote them by $q_{i}(B)$ and $q_{i}(D)$, in units of asset $B$ and $D$, respectively. The third bid is a tuple $Q_{i}=\{Q_{i}(B),Q_{i}(D)\}$, interpreted as a bid for the entire batch (i.e., for both orders) where $Q_{i}(B)$ is in units of asset $B$ and $Q_{i}(D)$ is in units of asset $D$.
@@
An important observation is that the feasibility constraint also has implications for the set of feasible mechanisms. Intuitively, a mechanism is feasible if, for any possible inventory, solvers can make non-trivial bids and be guaranteed to respect the feasibility constraint. The next definition formalizes this intuition.
@@
A mechanism $\Gamma$ is feasible if $\forall i\in\{1,2\}$,
$\exists\{q_{i}(B),q_{i}(D),Q_{i}(B),Q_{i}(D)\}>\{\underline{\beta},\underline{\delta},\underline{\beta},\underline{\delta}\}$
such that
@@ 3 Batch auction and simultaneous standard auctions
We can use the above framework to study the two dominant mechanisms for trade intent auctions: simultaneous standard auctions (as implemented by 1inch Fusion and UniswapX in the form of Dutch auctions) and batch auctions (as implemented by CoW Protocol). These mechanisms are quantitatively important: in October 2024, 1-inch fusion and UniswapX intermediated USD 831 million and USD 1.2 billion, respectively, while CoW Protocol intermediated USD 3.7 billion. These mechanisms are also relevant theoretical benchmarks because they are sub-cases of the above model: simultaneous standard auctions are equivalent to the above mechanism in which the batch bids are ignored, while batched auctions are equivalent to the above mechanism in which individual bids are ignored.
@@ 3.1 Simultaneous standard auctions
If there are two simultaneous standard auctions, then in equilibrium, solver 1 wins the first order and returns $\underline{\beta}$ units of asset $B$, while solver 2 wins the second order and returns $\underline{\delta}$ units of asset $D$ [^7] Note that, technically, there could be other equilibria because the least productive solver on each trade never wins. If the auction is, for example, in second price, then these other equilibria are even worse for the traders than that described in the Lemma. Also, these equilibria are not robust to perturbations of the type space. For example, if solver 1 can produce $\underline{\delta}+\epsilon$ for $\epsilon\rightarrow 0$ when it wins only trade 2, and solver 2 can produce $\underline{\beta}+\epsilon$ when it wins only trade 1, then there an infinitesimally small probability that solver 1 wins trade 2 and solver 2 wins trade 1. Then, the equilibrium described in the Lemma becomes unique.
@@ 3.2 Batch auction
To start, remember that feasibility prevents “second price” mechanisms that use the losing bid to determine the assets returned by the winning solver: even if the winning bid has a higher value, it may nonetheless deliver less of one asset than the second-highest bid. We, therefore, restrict our attention to a “first price” (or “pay your bid”) mechanism in which the winning solver returns the assets specified in its batched bid, which is the approach followed by CoW Protocol. [^9] In practice, the batch auction model implemented by CoW Protocol also has a subsidy to solvers, which we ignore for simplicity.
@@
The important observation is that, in the batch auction, the total value of the assets returned to the traders always exceeds that of the simultaneous individual-trade auctions. The reason is that batching forces the two solvers to compete: there is a single “strong” solver per trade and little competition if trades are auctioned off independently, but there are two “strong” solvers and high competition in the batch auction. At the same time, in the batch auction, there is an indeterminacy relative to what each trader receives. There are equilibria in which a trader receives less in a batch auction than what this trader would have received in the simultaneous individual-trade auctions. We summarize these observations in the following remark.
@@
Evaluated at the notional market price, traders collectively earn more in the batch auction than in the simultaneous standard auctions (c.f, Lemma 1). There are equilibria of the batch auction in which traders receive more than $\underline{\beta}$ and $\underline{\delta}$, and both traders are better off relative to the simultaneous standard auctions. However, there are also equilibria of the batch auction in which either trader 1 receives less than $\underline{\beta}$ or trader 2 receives less than $\underline{\delta}$; that is, there are equilibria in which one of the two traders is worse off relative to the simultaneous standard auctions.
@@
The goal of this section is to formalize a notion of fairness applicable to a generic combinatorial auction $\Gamma$. We do so by establishing a counterfactual mechanism against which to compare the outcome of $\Gamma$. The key observation is that such counterfactual is part of the mechanism $\Gamma$, because it depends on the bids and determines the equilibrium allocation and payments. That is, the reference for fairness must be generated endogenously as part of the mechanism.
@@
The first part of the above assumption imposes a mild form of monotonicity in the auction’s allocation. The second part is a form of independence of the auction’s payments whenever the individual bids determine the outcome. It is slightly more restrictive than the first part but still quite mild, especially in conjunction with the requirement that the mechanism is feasible. Remember that feasibility requires that a solver who does not win trade 1 does not return asset B and, similarly, a solver who does not win trade 2 does not return asset D. The above assumption then has implications exclusively in cases in which the individual bids determine the outcome and for the amount of asset B (asset D) returned by the winner of trade 1 (trade 2): these amounts should not depend on the batched bids.
@@ 5 Equilibrium of the fair combinatorial auction
The remainder of the section derives the equilibrium of different types of fair combinatorial auctions. For ease of comparison with the above benchmarks, we restrict our attention to mechanisms such that:
@@
    If multiple batched bids are fair, then the winner is the one with higher market value.
@@
Suppose that the individual, item-by-item auctions are second-price auctions. Formally, define the solver with the best batched bid as $W\equiv\mbox{argmax}_{i\in\{1,2\}}\left\{Q_{i}(B)+p_{DB}Q_{i}(D)\right\}$. The mechanism is: [^10] As already discussed in footnote 8, there can be some annoying technicalities related to ties — in this case, case a solver matches the opponent’s batched bid with its individual-trade bid. It turns out that, in this case, how these ties are broken does not matter. It will, however, matter in the next section when we consider the fair combinatorial auction with first-price individual auctions.
  • If $\forall i\in\{1,2\}$ we have $Q_{i}(B)\geq\min\{q_{1}(B),q_{2}(B)\}$ and $Q_{i}(D)\geq\min\{q_{1}(D),q_{2}(D)\}$, then the auction is equivalent to a batch auction: $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$ (and all other payments are zero).
  • If $\exists!i\in\{1,2\}$ such that $Q_{i}(B)\geq\min\{q_{1}(B),q_{2}(B)\}$ and $Q_{i}(D)\geq\min\{q_{1}(D),q_{2}(D)\}$, then the equilibrium is batching by solver $i$: $f(m)=\{i,i\}$, $b_{i}(m)=Q_{i}(B)$, $d_{i}(m)=Q_{i}(D)$, and all other payments are zero
  • Else, the best bid on each given order wins that order and pays the second-highest bid. Formally, denote the winner of order 1 by $w(B)\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(B)\}$, and the winner of order 2 by $w(D)\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(D)\}$. The mechanism in this case is: $f(m)=\{w(B),w(D)\}$, $b_{w(B)}(m)=\min\{q_{1}(B),q_{2}(B)\}$, $d_{w(D)}(m)=\min\{q_{1}(D),q_{2}(D)\}$ and all other transfers are zero.
Consider any strictly positive $Q_{1}(B),Q_{2}(B),Q_{1}(D),Q_{2}(D)$ that are an equilibrium of the batch auction discussed earlier. The key observation is that these bidding strategies are also an equilibrium of the fair combinatorial auction with individual bids in second price whenever $q_{1}(B)=q_{2}(B)\leq\min\{Q_{1}(B),Q_{2}(B)\}$ and $q_{1}(D)=q_{2}(D)\leq\min\{Q_{1}(D),Q_{2}(D)\}$. The reason is that no individual solver can manipulate the reference for fairness. That is, it is possible that, in equilibrium, $q_{1}(B)=q_{2}(B)\leq Q_{i}(B)<\underline{\beta}$ $\forall i\{1,2\}$; nonetheless, no solver can change its individual-trade bid and disqualify the opponent’s bid as unfair. This observation immediately leads to the following lemma.
@@
The set of equilibria of the fair combinatorial auction with second-price individual trade auctions (as in 2) is the set of equilibria of the batch auction.
Intuitively, when the individual-trade auctions are in second price, each individual solver has limited ability to manipulate the reference for fairness. Depending on the players’ bids, it is even possible that the outcome of the auction gives no asset to one player, and nonetheless, no solver can disqualify this outcome as unfair. As a consequence, this auction reduces to, again, the batch auction (c.f., Lemma 2).
@@
Again, call the solver with the best bid on trade 1 as $w(B)$, the solver with the best bid on trade 2 as $w(D)$, and the solver with the best batched bid by $W$ (see the beginning of the previous subsection for the definitions). It turns out that, unlike the previous case, here how ties are broken matters for the existence of the equilibrium. In particular, we need to distinguish between ties among bids from the same solver and ties among bids from different solvers.
Definition 4.
We say that the batched bid by solver $i\in\{1,2\}$ is fair whenever
$\begin{cases}Q_{i}(B)\geq q_{w(B)}&\mbox{ if }i\in w(B)\\ Q_{i}(B)>q_{w(B)}&\mbox{ otherwise }\end{cases}\hskip 28.45274pt\mbox{and}\hskip 28.45274pt\begin{cases}Q_{i}(D)\geq q_{w(D)}&\mbox{ if }i\in w(D)\\ Q_{i}(D)>q_{w(D)}&\mbox{ otherwise }\end{cases}$
Hence, a tie between a solver’s batched bid is “fair” if it ties with the same solver’s individual trade bid, but “unfair” if it ties with the opponent’s individual trade bid. The mechanism, in this case, is:
  • if both batched bids are fair, then $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$, and all other transfers are zero.
  • if only solver $i\in\{1,2\}$ batched bid is fair, then $f(m)=\{i,i\}$, $b_{i}(m)=Q_{i}(B)$, $d_{i}(m)=Q_{i}(D)$, and all other transfers are zero.
  • Otherwise, $f(m)=\{w(B),w(D)\}$, $b_{w(B)}(m)=q_{w(B)}(B)$, $d_{w(D)}(m)=q_{w(D)}(D)$ and all other transfers are zero.
Concerning the tie-breaking rule and why it is necessary for the existence of the equilibrium, note that if a solver wants the opponent’s batched bid to be disqualified as “unfair,” this solver needs to match the opponent’s batched bid with one of its individual trade bids. If the tie-breaking rule was different, then this solver would instead need to outbid the opponent’s bid by the smallest possible amount — which, however, is not defined because the bid space is continuous. Similarly, suppose that solver 1’s bid on trade 1 is sufficiently large to make solver 2’s batched bid “unfair”. Solver 1 will then want to set $Q_{1}(B)$ to the lowest possible level that does not disqualify its bid as “unfair”. Under our tie-breaking rule solver 1 can set $Q_{1}(B)=q_{1}(B)$, but under a different tie-breaking rule, this “lowest possible level” is not defined.
To derive the equilibrium, we start with a useful preliminary result.
@@
If $g>1$, then in every equilibrium of the game, it must be that $Q^{*}_{1}(D)=g\cdot\underline{\delta},\leavevmode\nobreak\ Q^{*}_{1}(B)\in(\underline{\beta},g\beta)$, and similarly $Q^{*}_{2}(B)=g\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)\in(\underline{\delta},g\cdot\delta)$.
If $g=1$, the equilibrium of the game is identical to the equilibrium of two simultaneous first-price auctions, as derived in Lemma 1. Without loss of generality, we can assume $Q^{*}_{1}(D)=g\cdot\underline{\delta},$ and $Q^{*}_{2}(B)=g\underline{\beta}$.
The lemma shows that, in its batched bid, each solver bids the maximum on the trade in which it has a disadvantage. This is the unique equilibrium when $g>1$ because there is a strictly positive probability that a batched bid wins, and is without loss of generality if $g=1$ because batched bids never win. The reason is that a solver’s batched bid is more likely to be considered unfair due to the amount of assets returned on the “weak” order (which is the opponent’s “strong” order). On the other component of the batched bid, each solver delivers more than the minimum amount but less than the maximum amount.
@@
$\begin{cases}(\beta-q_{1}(B))\mbox{pr}\left\{q_{2}(D)\geq g\underline{\delta}\right\}+\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)\geq g\underline{\beta}\\ \left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)<g\underline{\beta}\end{cases}$
where, again, $V\equiv g\underline{\beta}-p_{DB}\cdot g\underline{\delta}$. Note that, as long as $q_{1}(B)\geq g\underline{\beta}$, whether solver 1 wins one or both trades is independent of its bids, and solver 1 should set $q_{1}(B)=Q_{1}(B)=g\underline{\beta}$. A second observation is that if $q_{1}(B)<g\underline{\beta}$, then solver 1 payoff does not depend on $q_{1}(B)$ and, without loss of generality, we can assume $q_{1}(B)=\underline{\beta}$. In this case, $Q_{1}(B)$ is the best response to a first-price auction. We can therefore rewrite solver 1’s payoff as
$\begin{cases}(\beta-g\underline{\beta})\,\mathrm{pr}\{q_{2}(D)\geq g\underline{\delta}\}+(\beta\cdot g-g\underline{\beta})\,\mathrm{pr}\{q_{2}(D)<g\underline{\delta}\}&\text{if }q_{1}(B)=Q_{1}(B)=g\underline{\beta}\\ \\ \begin{aligned} &(\beta\cdot g-\tilde{Q}_{1}(B))\,\mathrm{pr}\{q_{2}(D)<g\underline{\delta}\}\\ &\quad\quad\times\,\mathrm{pr}\{\tilde{Q}_{1}(B)>p_{DB}Q_{2}(D)+V\mid q_{2}(D)<g\underline{\delta}\}\end{aligned}&\text{if }q_{1}(B)=\underline{\beta},Q_{1}(B)=\tilde{Q}_{1}(B)\end{cases}$ (4)
@@
$\tilde{Q}_{1}(B)\equiv\mbox{argmax}_{Q_{1}(B)}\left\{\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}\right\}.$
Following similar steps, by using Lemma 4 and eliminating dominated strategies, we can rewrite solver 2’s payoff as
$p_{DB}\cdot\begin{cases}(\delta-g\underline{\delta})\,\mathrm{pr}\{q_{1}(B)\geq g\underline{\beta}\}+(\delta\cdot g-g\underline{\delta})\,\mathrm{pr}\{q_{1}(B)<g\underline{\beta}\}&\text{if }q_{2}(D)=Q_{2}(D)=g\underline{\delta}\\ \\ \begin{aligned} &(\delta\cdot g-\tilde{Q}_{2}(D))\,\mathrm{pr}\{q_{1}(B)<g\underline{\beta}\}\\ &\quad\quad\times\,\mathrm{pr}\{p_{DB}Q_{2}(D)+V>Q_{1}(B)\mid q_{1}(B)<g\underline{\beta}\}\end{aligned}&\text{if }q_{1}(B)=\underline{\beta},Q_{1}(B)=\tilde{Q}_{1}(B)\end{cases}$ (5)
@@
$\tilde{Q}_{2}(D)\equiv\mbox{argmax}_{Q_{2}(D)}\left\{p_{DB}\left(\delta\cdot g-Q_{2}(D)\right)\mbox{pr}\left\{p_{DB}Q_{2}(D)+V>Q_{1}(B)|q_{1}(B)<g\underline{\beta}\right\}\right\}.$
The following Lemma shows that solver 1 always prefers $q_{1}(B)=g\underline{\beta}$ (i.e., kicking out the opponent’s batched bid) to $q_{1}(B)<g\underline{\beta}$ (i.e., not kicking out the opponent’s batched bid). However, $q_{1}(B)=g\underline{\beta}$ may not be feasible. Similarly, solver 2 sets $q_{2}(D)=g\underline{\delta}$ when feasible, and $q_{2}(D)=\underline{\delta}$ otherwise.
Lemma 5.
In every equilibrium of the game, solver 1 sets: $\begin{cases}q_{1}(B)=Q_{1}(B)=g\underline{\beta}&\mbox{ if }\beta\geq g\underline{\beta}\\ q_{1}(B)=\underline{\beta},\leavevmode\nobreak\ Q_{1}(B)=\tilde{Q}_{1}(B)&\mbox{ otherwise, }\end{cases}$ and solver 2 sets: $\begin{cases}q_{2}(D)=Q_{2}(D)=g\underline{\delta}&\mbox{ if }\delta\geq g\underline{\delta}\\ q_{2}(D)=\underline{\delta},\leavevmode\nobreak\ Q_{2}(D)=\tilde{Q}_{2}(D)&\mbox{ otherwise }\end{cases}$
@@
Figure 1: The two best responses (note: we only consider the bids not already pinned down by Lemma 4)
@@
The unique equilibrium of the fair combinatorial auction with first-price individual trade auctions (as in 1) is: